OpenStudy (tidbit233):

What is 3x-6y=6 in function notation?

OpenStudy (danjs):

y as a function of x y = f(x) = ... solve for y

OpenStudy (tidbit233):

Walk me through? Start with subtracting 3x?

OpenStudy (danjs):

sure, you can do that. You want to isolate 'y' by itself on one side. y = ..

OpenStudy (tidbit233):

y=(6-3x)/-6 ?

OpenStudy (danjs):

yeah you can simplify a bit more if you want, then write as y = f(x) = ...

OpenStudy (tidbit233):

How would it look simplified?

OpenStudy (tidbit233):

f(x)=3x + 1 ??

OpenStudy (danjs):

\[y = f(x) = \frac{ 6-3x }{ -6 }=\frac{ 6 }{ -6 }+\frac{ -3x }{ -6 }\] \[y=f(x)=-1 + \frac{ 1 }{ 2 }x\]

OpenStudy (tidbit233):

Thank you!